OpenStudy (warpedkitten):

Grant plans to evaporate enough water from 22 gallons of a 16% ammonia solution to make a 24% ammonia solution. Which equation can he use to find n, the number of gallons of water he should remove?

OpenStudy (mathstudent55):

22 gal of 16% ammonia has 0.16 * 22 = 3.52 gal of ammonia \(\dfrac{3.52}{22 - x} = 0.24\)